OpenStudy (anonymous):

How many different arrangements of the letters in the word PARALLEL are there? is it 40,320??

OpenStudy (lgbasallote):

PARALLEL has 8 letters A repeats twice L repeats thrice so the permutation becomes \[\frac{8!}{2!3!}\]

OpenStudy (anonymous):

13440?

OpenStudy (anonymous):

wait..

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

3,360?

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

\[3 \times 4 \times 5 \times 7 \times 8 = ??\]